OpenStudy (anonymous):

Given vectors u=(3,5) and v=4,-4), determine the quantity -4u--2v

jimthompson5910 (jim_thompson5910):

what is -4u?

OpenStudy (anonymous):

i'm not sure... if it helps, the question looks like

jimthompson5910 (jim_thompson5910):

-4u = -4(3,5) = (-4*3, -4*5) = (-12, -20) So -4u = (-12, -20) So what is -2v?

OpenStudy (anonymous):

(-8,8)

jimthompson5910 (jim_thompson5910):

good, so -4u--2v becomes (-12, -20) - (-8,8)

jimthompson5910 (jim_thompson5910):

then what's next?

OpenStudy (anonymous):

do you subtract -12--8 and -20-8?

jimthompson5910 (jim_thompson5910):

yes, you subtract the corresponding components

OpenStudy (anonymous):

so then its (-4,-28)

jimthompson5910 (jim_thompson5910):

bingo

OpenStudy (anonymous):

but the answers that i have to choose from are round numbers like -40 -256 -8 -64 88

jimthompson5910 (jim_thompson5910):

hmm one sec while I reread it

jimthompson5910 (jim_thompson5910):

maybe they want the dot product, I think i mistook that dot for a minus So this means that -4u dot -2v becomes (-12, -20) dot (-8,8) which turns into -12*(-8) + (-20)*(8) = 96 - 160 = -64

jimthompson5910 (jim_thompson5910):

so -4u dot -2v = -64

OpenStudy (anonymous):

oh, that makes more sense. i think that i did the same thing! no wonder it didn't work before.

jimthompson5910 (jim_thompson5910):

yeah strange how they used a - (well a small - in my opinion) instead of a dot

OpenStudy (anonymous):

yeah, i will have to keep an eye out for that next time. Thanks!

jimthompson5910 (jim_thompson5910):

you're welcome